OpenStudy (anonymous):

A building is being knocked down with a wrecking ball, which is a big metal sphere that swings on a 12-m-long cable. You are (unwisely!) standing directly beneath the point from which the wrecking ball is hung when you notice that the ball has just been released and is swinging directly toward you. How much time do you have to move out of the way?

OpenStudy (anonymous):

is that all the information you have?

OpenStudy (anonymous):

unfortunately

OpenStudy (anonymous):

i know that \[T= 2pi sqrt{m/k}\]

OpenStudy (anonymous):

\[T=2\pi \sqrt{m/k}\]

OpenStudy (anonymous):

so, \[T = 2\pi \sqrt{L/g} \] where L = Lenght g = gravity

OpenStudy (anonymous):

T- period?

OpenStudy (anonymous):

T- period in seconds i mean?

OpenStudy (anonymous):

yes, because you know that gravity is in m/s^2 so, period will be measured in seconds. You know the cable is 12m long and gravity is 9.81 solve for T (period) \[2\pi \sqrt{12/9.81} = 6.94922\]

OpenStudy (anonymous):

ok cool i got 6.95 as well

OpenStudy (anonymous):

but, The release through the first vertical point is ¼ of the total cycle, so you have ¼ of the period to get out of the way. \[6.95 * (1/4) = ?\]

OpenStudy (anonymous):

but that is in hertz correct?

OpenStudy (anonymous):

period is in seconds i believe

OpenStudy (anonymous):

ah ok. i ended up with 1.7 seconds

OpenStudy (anonymous):

sounds right to me.

OpenStudy (anonymous):

im still a little confused as to why we multiplied by 1/4

OpenStudy (anonymous):

Period is the time that the ball takes to complete one cycle, as you are standing right beneath the ball (first vertical point) that is 1/4 of the cycle

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

np (: